# Theory and Interpretability of Quantum Extreme Learning Machines: a Pauli-transfer matrix approach

### v1

@@ Appendix B Exact feature isolation criterion: null-space condition
Equation 38 is equivalent to requiring $\mathbf{w}R=\omega e_{r}^{\top}$ with some scalar weight $\omega$, i.e., $e_{r}^{\top}\in\mathbb{R}^{q}$ belongs to the row-space of the effective PTM $R\in\mathbb{R}^{{m}\times{q}}$. This, in turn, is equivalent to
@@
$\mathbf{w}=\omega\mathbf{e}_{r}^{\top}R^{+}$ (67)
@@
Figure 13: Null-space criterion (66): Number of isolatable features as a function of the number of observables (=rows of the PTM) m{m}, for (a) the TFIM at various evolution times $t$ and (b) a random unitary, for $n=4$ qubits ($d^{2}=256$). Panel (a) shows the total number of isolatable features, while panel (b) shows the number of isolatable 1-qubit features. In contrast to statistical decodability [cf. Appendix C], isolatability is an exact geometric criterion independent of the data.
Consider the case where $R\in\mathbb{R}^{{m}\times{q}}$ is dense and quasi-random, arising, e.g., for a Haar-random unitary $U$. For m<q{m}<{q}, the null space is necessarily nonzero and in random orientation. Accordingly, the probability that all vectors of $\mathrm{null}(R)$ have the same component vanish is essentially zero, such that no features are isolatable. Once ${m}\geq{q}$, $R$ has full column rank, hence $\mathrm{null}(R)=0$ and the null-space condition in Eq. 66 is trivially fulfilled, implying that all q{q} features become instantly isolatable for a random $R$. In this case one may define $W=R^{+}=R^{\top}$ and assign $\mathbf{w}=\Pi_{r}W$ with a $\Pi_{r}$ that projects onto the $r$-th coordinate. This $W$ fulfills the slightly more general condition of finding a matrix $W\in\mathbb{R}^{{m}\times{m}}$ such that $WR=[\mathbbm{1}_{{m}},0_{{q}-{m}}]$, thereby isolating all m{m} features.  [^10] Writing the PTM as $R=[R_{1},R_{2}]$ with $R_{2}\in\mathbb{R}^{{m}\times({q}-{m})}$, this requires in general $R_{2}=0$, which is not fulfilled for typical (random) unitaries except if symmetries are present and the basis is ordered accordingly (e.g., the parity symmetry of the TFIM would allow a diagonal block structure of $R$ with m=q/2{m}={q}/2).
For a Clifford unitary, which maps between elements of the Pauli group, $\mathrm{row}(R)$ simply consists of all $\mathbf{e}_{j}$, hence $\mathrm{rank}R$ features are isolatable even if $R$ is wide. For unitaries based on structured Hamiltonians or at small evolution times, one may expect Eq. 66 to be fulfilled for a certain set of features, provided m{m} is sufficiently large (but can still be ${m}\ll{q}$). This is confirmed numerically in Fig. 13, where the behavior of the feature isolation criterion (66) is illustrated for various unitaries as a function of the number of observables m{m} (rows of $R$).
@@ Appendix C Decodability scores and coefficient of determination
The decodability score $\gamma_{r}^{2}$ [Eq. 40] tells to which extent a feature can be geometrically reconstructed from the readout. It disregards statistical correlations and effects of noise, i.e., features with small singular values in the effective PTM $R$ could be considered decodable, although it would lead to huge weight norms. We present a refined $\mathcal{R}^{2}$-score below, but reserve a more detailed discussion of noise/regularization effects for a future work.
Consider the regression solution in Section IV.1, assume all quantities are centered and define the target alignment vector
$\mathbf{b}_{k}:=\mathbb{E}\!\left[\boldsymbol{\phi}(\mathbf{u})y_{k}(\mathbf{u})\right]\in\mathbb{R}^{q}.$ (68)
@@
$\mathbb{E}_{\mathbf{u}}\bigl[(y_{k}-f)^{2}\bigr]=\mathrm{Var}_{\mathbf{u}}(y_{k})-(R\mathbf{b}_{k})^{\top}\mathbf{w}^{*}.$ (69)
Plugging this into the definition in Eq. 43 gives the general expression for the coefficient of determination
$\mathcal{R}^{2}(k)=\frac{\mathbf{b}_{k}^{\top}R^{\top}(RCR^{\top})^{+}R\mathbf{b}_{k}}{\mathrm{Var}(y_{k})}.$ (70)
All quantities herein are to be averaged over the input distribution of $\mathbf{u}$. Notably, it is possible to have $\gamma_{k}^{2}=0$ (feature direction not present in row-space of $R$), yet still $\mathcal{R}^{2}(k)>0$ if $\phi_{r}$ is statistically correlated with other $\phi_{j}$ that are present. Hence, the above $\mathcal{R}^{2}$ generally measures predictability from correlations and not just isolatability of a feature.
The coefficient of determination $\mathcal{R}^{2}(k)$ can be related to the decodability score $\gamma_{r}^{2}$ [Eq. 40]: In the special where the targets are $y_{k}(\mathbf{u})=\phi_{k}(\mathbf{u})$, and under the additional assumption of orthonormality,
$\mathbb{E}[y_{k}y_{k^{\prime}}]=\delta_{kk^{\prime}}\implies C=\mathbbm{1}_{q},\quad\mathbf{b}_{k}=\mathbf{e}_{k},\quad\mathrm{Var}(y_{k})=1$ (71)
Eq. 70 reduces to the decodability score
$\mathcal{R}^{2}(k)=\mathbf{b}_{k}^{\top}R^{\top}(RR^{\top})^{+}R\mathbf{b}_{k}=\mathbf{e}_{k}^{\top}(R^{+}R)\mathbf{e}_{k}=\gamma_{k}^{2}.$ (72)
In the generic non-orthogonal case, the exact relation breaks down and $\mathcal{R}^{2}$ depends on the full projector $R^{+}R$ and not just on its diagonal elements. Then, $\mathcal{R}^{2}$ measures the extent to which the target lies in the decodable subspace.
@@ D.1 Unitary channels
Consider a separable reservoir Hamiltonian $H_{R}=\sum_{j=1}^{n}h_{R}^{(j)}$, such that the reservoir unitary factorizes as $U(t)=\bigotimes_{j=1}^{n}U_{j}(t)$. Such a structure emerges, e.g., in the limit of large transverse field $h/J\gg 1$ of the TFIM (5) This implies a Kronecker-product structure of the unitary PTM (73) (and similarly for any other separable quantum channel),
@@
$U_{k}(t)^{\dagger}XU_{k}(t)=\cos(2t)\,X+\sin(2t)\,Y,\qquad U_{k}(t)^{\dagger}YU_{k}(t)=\cos(2t)\,Y-\sin(2t)\,X,$ (78)
so that $\phi_{y}^{{}^{\prime}(k)}(u_{k})=-\sin(2t)\,\phi_{x}^{(k)}(u_{k})$ and, correspondingly, $\phi^{\prime}_{\ldots y_{k}\ldots}(\mathbf{x})=-\sin(2t)\,\phi_{\ldots x_{k}\ldots}(\mathbf{x})$, while $Z$-features remain invariant. This shows how unitaries act to make specific features accessible to the readout when the latter uses a restricted set of observables.
@@ D.2 General quantum channels: affine action on the Bloch part
where $\mathbf{t}=\mathbf{0}$ for unital channels and $\Omega$ is a contracting map (singular values $\leq 1$). Such a map can improve numerical conditioning by damping high-weight Pauli components that are often noise sensitive. Non-unitality produces a translation $\mathbf{t}$ of the Bloch part, e.g., a maximally mixed state would be shifted away from the center of the Bloch sphere. This becomes a learnable bias once combined with the identity component of an observable.
@@
    For an $n$-qubit isotropic depolarizing channel on dimension $d=2^{n}$, $T_{\mathrm{dep}}^{(n)}(\lambda)=\mathrm{diag}\big(1,\lambda I_{d^{2}-1}\big)$.
@@
Modeling $n$-qubit noise as tensor products of single-qubit channels $\mathcal{E}^{(n)}=\bigotimes_{i=1}^{n}\mathcal{E}^{(i)}$ is often used as a baseline in QRC models [115, 18, 21, 115, 46]. Since the PTMs (81) and (82) are Pauli-diagonal, they do not lead to Pauli spreading and hence do not enlarge the accessible feature subspace. Amplitude damping [Eq. 83] produces Pauli-I/Z-mixing, which causes operator weight decay.
@@
$\mathcal{D}^{(n)}(P)=\lambda^{\nu(P)}P$ (84)
Accordingly, high-weight Pauli strings are exponentially suppressed by near-diagonal noise channels. In conjunction with unitary-induced operator spreading, this can increase information loss under finite sampling. On the other hand, amplitude damping noise was observed to improve QRC performance as the associated superoperator spreads the state’s probability weight into more Pauli directions, thereby enriching the feature map [18]. This has been related to the ’coherence influx’ present for amplitude damping but not depolarization [43].

### v2

@@ Appendix B Exact feature isolation criterion: null-space condition
Equation 39 is equivalent to requiring $\mathbf{w}^{\top}R=\omega e_{r}^{\top}$ with some scalar weight $\omega$, i.e., $e_{r}^{\top}\in\mathbb{R}^{q}$ belongs to the row-space of the effective PTM $R\in\mathbb{R}^{{m}\times{q}}$. This, in turn, is equivalent to
@@
$\mathbf{w}^{\top}=\omega\mathbf{e}_{r}^{\top}R^{+}$ (63)
@@
Figure 14: Null-space criterion (62): Number of isolatable features as a function of the number of observables (=rows of the PTM) m{m}, for (a) the TFIM at various evolution times $t$ and (b) a random unitary, for $n=4$ qubits ($d^{2}=256$). Panel (a) shows the total number of isolatable features, while panel (b) shows the number of isolatable 1-qubit features. In contrast to statistical decodability (cf. Appendix C), isolatability is an exact geometric criterion independent of the data.
Consider the case where $R\in\mathbb{R}^{{m}\times{q}}$ is dense and quasi-random, arising, e.g., for a Haar-random unitary $U$. For m<q{m}<{q}, the null space is necessarily nonzero and in random orientation. Accordingly, the probability that the same component vanishes for all vectors of $\mathrm{null}(R)$ is essentially zero, such that no features are isolatable. Once ${m}\geq{q}$, $R$ has full column rank, hence $\mathrm{null}(R)=0$ and the null-space condition in Eq. 62 is trivially fulfilled, implying that all q{q} features become instantly isolatable for a random $R$. In this case one may define $W=R^{+}=(R^{\top}R)^{-1}R^{\top}$ and assign $\mathbf{w}^{\top}=\mathbf{e}_{r}^{\top}W$.
For a Clifford unitary, which maps between elements of the Pauli group, $\mathrm{row}(R)$ simply consists of a subset of signed standard basis vectors, hence $\mathrm{rank}(R)$ features are isolatable even if $R$ is wide. For unitaries based on structured Hamiltonians or at small evolution times, one may expect Eq. 62 to be fulfilled for a certain set of features, provided m{m} is sufficiently large (but can still be ${m}\ll{q}$). This is confirmed numerically in Fig. 14, where the behavior of the feature isolation criterion (62) is illustrated for various unitaries as a function of the number of observables m{m} (rows of $R$).
@@ Appendix C Decodability scores and coefficient of determination
The decodability score $\gamma_{r}^{2}$ [Eq. 41] tells to which extent a feature can be geometrically reconstructed from the readout. It disregards statistical correlations and effects of noise, i.e., features with small singular values in the effective PTM $R$ could be considered decodable, although it would lead to huge weight norms. A straightforward way to limit noise amplification is to replace the pseudoinverse by its Tikhonov regularized form
$\gamma_{r,\lambda}^{2}=(R_{\lambda}^{+}R)_{rr}=\left[(R^{\top}R+\lambda\mathbbm{1})^{-1}R^{\top}R\right]_{rr}=\sum_{i=1}^{s}\frac{\sigma_{i}^{2}}{\sigma_{i}^{2}+\lambda}v_{ri}^{2},\qquad\gamma_{r,\lambda\to 0}^{2}\to\gamma_{r}^{2}$ (64)
where $\sigma_{i=1,\ldots,s}$ are the singular values of $R$ and $\mathbf{v}_{i}$ the right-singular vectors. Directions with $\sigma_{i}^{2}\ll\lambda$ contribute little to $\gamma_{r,\lambda}^{2}$ and are thus considered undecodable.
To account for data correlations within a refined $\mathcal{R}^{2}$-score, consider the regression solution in Section IV.1, assume all quantities are centered, and define the target alignment vector
$\mathbf{b}_{k}:=\mathbb{E}\!\left[\boldsymbol{\upphi}(\mathbf{u})y_{k}(\mathbf{u})\right]\in\mathbb{R}^{q}.$ (65)
@@
$\mathbb{E}_{\mathbf{u}}\bigl[(y_{k}-f)^{2}\bigr]=\mathrm{Var}_{\mathbf{u}}(y_{k})-(R\mathbf{b}_{k})^{\top}\hat{\mathbf{w}}$. (66)
Plugging this into the definition in Eq. 44 gives the expression for the coefficient of determination (in the unregularized case $\lambda=0$)
$\mathcal{R}^{2}[y_{k}]=\frac{\mathbf{b}_{k}^{\top}R^{\top}(RCR^{\top})^{+}R\mathbf{b}_{k}}{\mathrm{Var}(y_{k})},$ (67)
where $C=\mathbb{E}[\boldsymbol{\upphi}\boldsymbol{\upphi}^{\top}]$ is the feature covariance matrix. All quantities herein are to be averaged over the input distribution of $\mathbf{u}$.
In the special where the targets are the Pauli features, $y_{r}(\mathbf{u})=\phi_{r}(\mathbf{u})$, and under the additional assumption of orthonormality (whitened features),
$\mathbb{E}[y_{r}y_{r^{\prime}}]=\delta_{rr^{\prime}}\implies C=\mathbbm{1}_{q},\quad\mathbf{b}_{r}=\mathbf{e}_{r},\quad\mathrm{Var}(y_{r})=1,$ (68)
the coefficient of determination $\mathcal{R}^{2}[y_{r}]$ in Eq. 67 reduces to the decodability score
$\mathcal{R}^{2}[y_{r}]=\mathbf{b}_{r}^{\top}R^{\top}(RR^{\top})^{+}R\mathbf{b}_{r}=\mathbf{e}_{r}^{\top}(R^{+}R)\mathbf{e}_{r}=\gamma_{r}^{2}.$ (69)
In the generic non-orthogonal case, this simple relation breaks down and $\mathcal{R}^{2}$ depends also on the covariance $C$ and target-feature alignment. Then, $\mathcal{R}^{2}$ measures the extent to which the target lies in the decodable subspace. As a consequence, it is then possible to have $\gamma_{r}^{2}=0$ (feature direction not present in row-space of $R$), yet still $\mathcal{R}^{2}[y_{r}]>0$ if $\phi_{r}$ is statistically correlated with other $\phi_{j}$ that are present. Hence, $\mathcal{R}^{2}$ in Eq. 67 generally measures predictability from correlations and not just isolatability of a feature.
@@ D.1 Unitary channels
Consider a separable reservoir Hamiltonian $H_{R}=\sum_{j=1}^{n}h_{R}^{(j)}$, such that the reservoir unitary factorizes as $U(t)=\bigotimes_{j=1}^{n}U_{j}(t)$. Such a structure emerges, e.g., for a TFIM with $h/J\to\infty$ [Eq. 5]. This implies a Kronecker-product structure of the unitary PTM (70) (and similarly for any other product quantum channel),
@@
$U_{k}(t)^{\dagger}XU_{k}(t)=\cos(2t)\,X-\sin(2t)\,Y,\qquad U_{k}(t)^{\dagger}YU_{k}(t)=\cos(2t)\,Y+\sin(2t)\,X,$ (75)
so that $\phi_{y}^{{}^{\prime}(k)}(u_{k})=\sin(2t)\,\phi_{x}^{(k)}(u_{k})$ and, correspondingly, $\phi^{\prime}_{\ldots y_{k}\ldots}(\mathbf{x})=\sin(2t)\,\phi_{\ldots x_{k}\ldots}(\mathbf{x})$, while $Z$-features remain invariant. This shows how unitaries act to make specific features accessible to the readout when the latter uses a restricted set of observables.
@@ D.2 General quantum channels: affine action on the Bloch part
where $\mathbf{t}=\mathbf{0}$ for unital channels. For a single qubit, $\Omega$ is a contracting map (singular values $\leq 1$). Such a map can improve numerical robustness by damping high-weight Pauli components that are often noise sensitive. Non-unitality produces a translation $\mathbf{t}$ of the Bloch part, e.g., a maximally mixed state would be shifted away from the center of the Bloch sphere.
@@
    For a global isotropic depolarizing channel on dimension $d=2^{n}$, $T_{\mathrm{dep}}^{(n)}(\lambda)=\mathrm{diag}\big(1,\lambda I_{d^{2}-1}\big)$.
@@
Modeling $n$-qubit noise as tensor products of single-qubit channels $\mathcal{E}^{(n)}=\bigotimes_{i=1}^{n}\mathcal{E}^{(i)}$ is often used as a baseline in QRC models [129, 19, 22, 49]. Since the PTMs (78) and (79) are Pauli-diagonal, they do not lead to Pauli spreading and hence do not enlarge the accessible feature subspace. Amplitude damping [Eq. 80] produces Pauli-I/Z-mixing, which causes operator weight decay.
@@
$\mathcal{D}^{(n)}(P)=\lambda^{\nu(P)}P.$ (81)
Accordingly, high-weight Pauli strings are exponentially suppressed by near-diagonal noise channels. In conjunction with unitary-induced operator spreading, this can increase information loss under finite sampling. On the other hand, amplitude damping noise is often observed to improve QRC performance as the associated superoperator spreads the state’s probability weight into more Pauli directions, thereby enriching the feature map [19]. This has been related to the ‘coherence influx’ present for amplitude damping but not depolarization [46].